# This is the extended version of the conference version at FoSSaCS 2022 [BFPZ22] [a] [a] [a]

### v1

@@ 3. Strict and global probability-raising causes
Our contribution starts by providing novel formal definitions for cause-effect relations in MDPs which rely on the probability-raising (PR) principle as stated by (C1) and (C2) in the introduction. Here, we focus on the case where both causes and effects are state properties, i.e., sets of states.
In the sequel, let $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ be an MDP and $\mathsf{Eff}\subseteq S\setminus\{\mathsf{init}\}$ a nonempty set of terminal states. As the effect set is fixed, the assumption that all effect states are terminal contributes to the temporal priority as stated by (C2). We may also assume that every state $s\in S$ is reachable from $\mathsf{init}$.
We consider two variants of the probability-raising condition: the global setting treats the set $\mathsf{Cause}$ as a unit, while the strict view requires the (C1) for all states in $\mathsf{Cause}$ individually.
@@ 3.1. Examples and simple properties of probability-raising causes
Figure 1. A MC allowing for non-strict GPR causes
$\mathsf{init}$$\mathsf{eff}$11
Figure 2. A MC in which no PR cause exists
@@
Let $\mathcal{M}$ be the MDP in Figure 4, where the notation is similar to Example 3.1 with the addition of actions $\alpha,\beta$ and $\gamma$. Let $\mathsf{Cause}=\{c\}$ and $\mathsf{Eff}=\{\mathsf{eff}\}$. Only state $s$ has a nondeterministic choice. $\mathsf{Cause}$ is not an PR cause. To see this, regard the finite-memory deterministic scheduler $\mathfrak{T}$ that schedules $\beta$ only for the first visit of $s$ and $\alpha$ for the second visit of $s$. Then:
@@
[…]
Figure 3. MDP $\mathcal{M}$ from Remark 6
[…]
Figure 4. MDP $\mathcal{M}$ from Remark 7
@@
Remark 9 (Action causality and other forms of PR causality).
@@
Suppose, the task is to check whether taking action $\alpha$ in state $s$ raises the effect probabilities compared to never scheduling $\alpha$ in state $s$. Let $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ be copies of $\mathcal{M}$ with the following modifications: In $\mathcal{M}_{0}$, the only enabled action of state $s$ is $\alpha$, while in $\mathcal{M}_{1}$ the enabled actions of state $s$ are the elements of $\mathit{Act}_{\mathcal{M}}(s)\setminus\{\alpha\}$. Let now $\mathcal{N}$ be the MDP whose initial state has a single enabled action and moves with probability $1/2$ to $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$. Then, action $\alpha$ raises the effect probability in $\mathcal{M}$ iff the copy of the state $s$ in $\mathcal{M}_{0}$ constitutes an SPR cause for both copies of the effect in $\mathcal{N}$. This idea can be generalized to check whether scheduler classes satisfying a regular constraint have higher effect probability compared to all other schedulers. Here, we can deal with an MDP $\mathcal{N}$ as above where $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ are defined as the synchronous product of deterministic finite automata and $\mathcal{M}$.
@@
To demonstrate this consider the MDP $\mathcal{M}$ from Figure 6. We are interested in reaching the effect state $\mathsf{eff}$. For the action $\delta$ in state $s_{1}$ the corresponding construction is depicted in Figure 6. The state $s_{1,0}$ as the copy of $s_{1}$ in $\mathcal{M}_{0}$ satisfies (SPR), since for any scheduler $\mathfrak{S}$ reaching $s_{1,0}$ we have
$\mathrm{Pr}_{\mathcal{N},s_{1,0}}^{\mathfrak{S}}(\lozenge\{\mathsf{eff}_{0},\mathsf{eff}_{1}\})=1/4>107/800\geq\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\lozenge\{\mathsf{eff}_{0},\mathsf{eff}_{1}\})$,
[…]
@@ 4. Checking the existence of PR causes and the PR conditions
We now turn to algorithms for checking whether a given set $\mathsf{Cause}$ is an SPR or GPR cause for $\mathsf{Eff}$. Since the minimality condition (for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$) of PR causes is verifiable by standard model checking techniques in polynomial time, we concentrate on checking the probability-raising conditions (S) and (G). For Markov chains, both (SPR) and (GPR) can be checked in polynomial time by computing the corresponding probabilities. So, the interesting case is checking the PR conditions in MDPs. In case of SPR causality, this is closely related to the existence of PR causes and solvable in polynomial time (Section 4.1), while checking the GPR condition is more complex and polynomially reducible to (the non-solvability of) a quadratic constraint system (Section 4.2).
We start by stating that for (S) and (G), it suffices to consider schedulers minimizing the probability to reach an effect state from every cause state. For this we transform the MDP in question according to the cause candidate in question.
@@
We fix a state $c\in\mathsf{Cause}$. Recall also that we assume the states in $\mathsf{Eff}$ to be terminal. Let $\psi_{c}=(\neg\mathsf{Cause})\,\mathrm{U}\,c$, $w_{c}=\mathrm{Pr}_{\mathcal{M},c}^{\min}(\Diamond\mathsf{Eff})$ and let $\Upsilon_{c}$ denote the set of all schedulers $\mathfrak{U}$ for $\mathcal{M}$ such that
@@
$\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond\mathsf{Cause})>0$
and
@@
It now suffices to show that for each scheduler $\mathfrak{S}\in\Sigma_{>0}$ there exists a scheduler $\mathfrak{S}^{\prime}\in\Sigma_{>0,\min}$ such that if (GPR) holds $\mathfrak{S}^{\prime}$ then (GPR) holds for $\mathfrak{S}$. So, let $\mathfrak{S}\in\Sigma_{>0}$.
@@ 4.1. Checking the strict probability-raising condition and the existence of causes
First, we show the soundness of Algorithm 4.1. By the virtue of Lemma 10 it suffices to show that Algorithm 4.1 returns the correct answers “yes” or “no” when the task is to check whether the singleton $\mathsf{Cause}=\{c\}$ is an SPR cause in $\mathcal{N}=\mathcal{M}_{[c]}$. Recall the notation $q_{s}=\mathrm{Pr}^{\max}_{\mathcal{M}_{[c]},s}(\Diamond\mathsf{Eff})$. We abbreviate $q=q_{\mathsf{init}}$. Note that $(\neg\mathsf{Cause})\,\mathrm{U}\,c$ is equivalent to $\Diamond c$.
@@
Obviously, statement (c) implies statements (a) and (b). The implication “(a) $\Longrightarrow$ (b)” follows from Lemma 5. We now turn to the proof of “(b) $\Longrightarrow$ (c)”. For this, we assume that we are given a GPR cause $\mathsf{Cause}$ for $\mathsf{Eff}$ in $\mathcal{M}$. For $c\in\mathsf{Cause}$, let $w_{c}=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$. Pick a state $c_{0}\in\mathsf{Cause}$ such that $w_{c_{0}}=\max\{w_{c}:c\in\mathsf{Cause}\}$. For every scheduler $\mathfrak{S}$ for $\mathcal{M}$ that minimizes the effect probability whenever it visits a state in $\mathsf{Cause}$, and visits $\mathsf{Cause}$ with positive probability, the conditional probability $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff}|\Diamond\mathsf{Cause})$ is a weighted average of the values $w_{c}$, $c\in\mathsf{Cause}$, and thus bounded by $w_{c_{0}}$. Using Lemma 10 it is now easy to see that $\{c_{0}\}$ is both an SPR and a GPR cause for $\mathsf{Eff}$.
@@ 4.2. Checking the global probability-raising condition
The remainder of this section is concerned with the proofs of Theorem 17 and Theorem 18. We suppose that $\mathsf{Cause}$ satisfies for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$.
@@
The first step is a polynomial-time model transformation which permits to make the following assumptions when checking the GPR condition of $\mathsf{Cause}$ for $\mathsf{Eff}$.
@@
For every state $c\in\mathsf{Cause}$, there is only a single enabled action, say $\mathit{Act}(c)=\{\gamma\}$, and there exists $w_{c}\in[0,1]\cap\mathbb{Q}$ such that $P(c,\gamma,\mathsf{eff}_{\mathsf{cov}})=w_{c}$ and $P(c,\gamma,\mathsf{noeff}_{\mathsf{fp}})=1{-}w_{c}$ where $\mathsf{noeff}_{\mathsf{fp}}$ is a terminal non-effect state and $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{eff}_{\mathsf{cov}}$ are only accessible via the $\gamma$-transition from the states $c\in\mathsf{Cause}$.
@@
The terminal states $\mathsf{eff}_{\mathsf{unc}}$, $\mathsf{eff}_{\mathsf{cov}}$, $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{noeff}_{\mathsf{tn}}$ are supposed to be pairwise distinct. $\mathcal{M}$ can have further terminal states representing true negatives. These could be identified with $\mathsf{noeff}_{\mathsf{tn}}$, but this is irrelevant for our purposes.
@@
We justify the assumptions as we can transform $\mathcal{M}$ into a new MDP of the same asymptotic size satisfying the above assumptions. Thanks to Lemma 10, we may suppose that $\mathcal{M}=\mathcal{M}_{[\mathsf{Cause}]}$ (see Notation 4) without changing the satisfaction of (G). We then may rename the effect state $\mathsf{eff}$ and the non-effect state $\mathsf{noeff}$ reachable from $\mathsf{Cause}$ into $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$, respectively. Furthermore, we collapse all other effect states into a single state $\mathsf{eff}_{\mathsf{unc}}$ and all true negative states into $\mathsf{noeff}_{\mathsf{tn}}$. Similarly, by renaming and possibly duplicating terminal states we also suppose that $\mathsf{noeff}_{\mathsf{fp}}$ has no other incoming transitions than the $\gamma$-transitions from the states in $\mathsf{Cause}$. This ensures (A1) and (A2). For (A3) consider the set $T$ of terminal states in the MDP obtained so far. We remove all end components by switching to the MEC-quotient [dA97], i.e., we collapse all states that belong to the same MEC $\mathcal{E}$ into a single state $s_{\mathcal{E}}$ while ignoring the actions inside $\mathcal{E}$. Additionally, we add a fresh $\tau$-transition from the states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ (i.e., $P(s_{\mathcal{E}},\tau,\mathsf{noeff}_{\mathsf{tn}})=1$). The $\tau$-transitions from states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ mimic cases where schedulers of the original MDP eventually enter an end component and stay there forever with positive probability.
@@
By Lemma 3, there is a scheduler $\mathfrak{T}$ for $\mathcal{N}$ for each scheduler $\mathfrak{S}$ for $\mathcal{M}^{\prime}$ such that each terminal state is reached with the same probability under $\mathfrak{T}$ in $\mathcal{N}$ and under $\mathfrak{S}$ in $\mathcal{M}^{\prime}$. The state $\mathsf{eff}_{\mathsf{cov}}$ is also present in $\mathcal{M}_{[\mathsf{Cause}]}$ under the name $\mathsf{eff}$ and reached with the same probability as in $\mathcal{M}^{\prime}$ when $\mathfrak{S}$ is considered as a scheduler for $\mathcal{M}_{[\mathsf{Cause}]}$. The state $\mathsf{eff}$ is furthermore reached in $\mathcal{M}_{[\mathsf{Cause}]}$ if and only if $\Diamond\mathsf{Cause}\land\Diamond\mathsf{Eff}$ is satisfied along a run. The set of terminal states in $\mathsf{Eff}$ is obtained from the set $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$ by collapsing states. As a scheduler $\mathfrak{S}$ can be viewed as a scheduler for both MDPs and these MDPs agree except for the terminal states, the first equality follows as well. As the probability to reach $\mathsf{Cause}$ is the sum of the probabilities to reach the terminal states $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$ in $\mathcal{N}$ and $\mathcal{M}^{\prime}$ and as these states are only renamed in the transition from $\mathcal{M}_{[\mathsf{Cause}]}$ to $\mathcal{M}^{\prime}$, the claim follows.
@@
Note, however, that the transformation changes the memory-requirements of schedulers witnessing that $\mathsf{Cause}$ is not a GPR cause for $\mathsf{Eff}$. We will address the memory requirements in the original MDP later. With assumptions (A1)-(A3), (G) can be reformulated as follows:
@@
With assumptions (A1)-(A3), a terminal state of $\mathcal{M}$ is reached almost surely under any scheduler after finitely many steps in expectation. Given a scheduler $\mathfrak{S}$ for $\mathcal{M}$, the expected frequencies (i.e., expected number of occurrences in maximal paths) of state action-pairs $(s,\alpha)$, states $s\in S$ and state-sets $T\subseteq S$ under $\mathfrak{S}$ are defined by:
@@
This now puts us in the position to prove Theorem 17.
@@
Let $B_{2}$ be an open ball in $\mathbb{R}^{r}$ such that $h(B_{2})\subseteq B_{1}$ and such that $h(B_{2})$ contains $(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. We claim that $g\circ h\colon B_{2}\to\mathbb{R}$ has an isolated local maximum at $z^{\ast}\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}h^{-1}(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. It is clear that $g\circ h$ has local maximum since $g$ has a local maximum at $(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. Suppose now, that $g\circ h$ does not have an isolated local maximum at $h^{-1}(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. As $h$ is an affine map and the graph of $g$ is the solution to a quadratic equation, this is only possible if there is a direction $d\in\mathbb{R}^{r}\setminus\{0\}$ such that

### v2

@@ 3. Strict and global probability-raising causes
Our contribution starts by providing novel formal definitions for cause-effect relations in MDPs which rely on the probability-raising (PR) principle $P(E\mid C)>P(C)$ stated by (C1) and temporal priority of causes as stated by (C2) in the introduction. Here, we focus on the case where both causes and effects are state properties, i.e., sets of states.
In the sequel, let $\mathcal{M}=(S,\mathit{Act},P,\mathsf{init})$ be an MDP and $\mathsf{Eff}\subseteq S\setminus\{\mathsf{init}\}$ a nonempty set of terminal states. As the effect set is fixed, the assumption that all effect states are terminal contributes to the temporal priority (C2). We may also assume that every state $s\in S$ is reachable from $\mathsf{init}$.
We consider two variants of the probability-raising condition: the global setting treats the set $\mathsf{Cause}$ as a unit, while the strict view requires (C1) for all states in $\mathsf{Cause}$ individually.
@@ 3.1. Examples and simple properties of probability-raising causes
Figure 1. A MC allowing for non-strict GPR causes
$\mathsf{init}$$\mathsf{eff}$11
Figure 2. A MC in which no PR cause exists
@@
Let $\mathcal{M}$ be the MDP in Figure 4, where the notation is similar to Example 2 with the addition of actions $\alpha,\beta$ and $\gamma$. Let $\mathsf{Cause}=\{c\}$ and $\mathsf{Eff}=\{\mathsf{eff}\}$. Only state $s$ has a nondeterministic choice. $\mathsf{Cause}$ is not an PR cause. To see this, regard the finite-memory deterministic scheduler $\mathfrak{T}$ that schedules $\beta$ only for the first visit of $s$ and $\alpha$ for the second visit of $s$. Then:
@@
[…]
Figure 3. MDP $\mathcal{M}$ from Remark 6
[…]
Figure 4. MDP $\mathcal{M}$ from Remark 7
@@
Remark 9 (Action PR causality).
@@
Suppose, the task is to check whether taking action $\alpha$ in state $s$ raises the effect probabilities compared to never scheduling $\alpha$ in state $s$. This form of action causality was discussed in an example in [DFT20]. We argue that we can deal with this kind of causality to. For this we assume there are no cycles in $\mathcal{M}$ containing $s$. Let $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ be copies of $\mathcal{M}$ with the following modifications: In $\mathcal{M}_{0}$, the only enabled action of state $s$ is $\alpha$, while in $\mathcal{M}_{1}$ the enabled actions of state $s$ are the elements of $\mathit{Act}_{\mathcal{M}}(s)\setminus\{\alpha\}$.
The action $\alpha$ raises the effect probability in $\mathcal{M}$ if and only if for all scheduler $\mathfrak{S}$ of $\mathcal{N}$ the copy of $s$ in $\mathcal{M}_{0}$ satisfies (SPR) in $\mathcal{N}$. This idea can be generalized to check whether scheduler classes satisfying a regular constraint have higher effect probability compared to all other schedulers. In this case, we can deal with an MDP $\mathcal{N}$ as above where $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$ are defined as the synchronous product of deterministic finite automata and $\mathcal{M}$.
To demonstrate this consider the MDP $\mathcal{M}$ from Figure 6. We are interested whether taking $\alpha$ in $s$ raises the probability to reach the effect state $\mathsf{eff}$. The constructed MDP $\mathcal{N}$ with two adapted copies of $\mathcal{M}$ is depicted in Figure 6. For all scheduler $\mathfrak{S}$ of $\mathcal{N}$ the state $s_{0}$ satisfies (SPR) by
$\mathrm{Pr}_{\mathcal{N}}^{\mathfrak{S}}(\lozenge\mathsf{eff}\mid\lozenge s_{0})=1/4>1/8=\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{N}}(\lozenge\mathsf{eff})$,
[…]
@@ 4. Checking the existence of PR causes and the PR conditions
We now turn to algorithms for checking whether a given set $\mathsf{Cause}$ is an SPR or GPR cause for $\mathsf{Eff}$ in $\mathcal{M}$. Since the minimality condition (for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$) of PR causes is verifiable by standard model checking techniques in polynomial time, we concentrate on checking the probability-raising conditions (S) and (G). For Markov chains, both (SPR) and (GPR) can be checked in polynomial time by computing the corresponding probabilities. So, the interesting case is checking the PR conditions in MDPs. In case of SPR causality this is closely related to the existence of PR causes and solvable in polynomial time (Section 4.1), while checking the GPR condition is more complex and polynomially reducible to (the non-solvability of) a quadratic constraint system (Section 4.2).
We start by stating that for (S) and (G), it suffices to consider kind of worst-case scheduler, which are minimizing the probability to reach an effect state from every cause state. For this we transform the MDP in question according to the cause candidate in question.
@@
{exa}
[…]
Figure 7. Example MDP $\mathcal{M}$
[…]
Figure 8. Transformed MDP $\mathcal{M}_{[c]}$
@@
We show that $\mathsf{Cause}$ is an SPR cause in $\mathcal{M}$ by showing (S) for all states in $\mathsf{Cause}$. Thus, we fix a state $c\in\mathsf{Cause}$. Recall also that we assume the states in $\mathsf{Eff}$ to be terminal. Let $\psi_{c}=(\neg\mathsf{Cause})\,\mathrm{U}\,c$, $w_{c}=\mathrm{Pr}_{\mathcal{M},c}^{\min}(\Diamond\mathsf{Eff})$ and let $\Upsilon_{c}$ denote the set of all schedulers $\mathfrak{U}$ for $\mathcal{M}$ such that
@@
$\mathrm{Pr}_{\mathcal{M}}^{\mathfrak{S}}(\Diamond\mathsf{Cause})>0\qquad\text{and}\qquad\mathrm{Pr}^{\mathit{res}(\mathfrak{S},\pi)}_{\mathcal{M},c}(\Diamond\mathsf{Eff})\ \ =\ \ \mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$
@@
It now suffices to show that for each scheduler $\mathfrak{S}\in\Sigma_{>0}$ there exists a scheduler $\mathfrak{S}^{\prime}\in\Sigma_{>0,\min}$ such that if (GPR) holds for $\mathfrak{S}^{\prime}$ then (GPR) holds for $\mathfrak{S}$. So, let $\mathfrak{S}\in\Sigma_{>0}$.
@@ 4.1. Checking the strict probability-raising condition and the existence of causes
As the construction of the MDP $\mathcal{M}_{[c]}$ suggests, the two values compared by the algorithm are instances of worst-case scheduler. On one hand, the probability to reach $\mathsf{Eff}$ starting in $c$ is minimized, while it is maximized if $c$ was not seen yet. If in such a scenario we have case 1. $q_{\mathsf{init}}<w_{c}$ then $c$ obviously satisfies (SPR). In the case 2. $q_{\mathsf{init}}>w_{c}$ we can build a scheduler which refuses (SPR) for $c$. Lastly, in the corner case 3. $q_{\mathsf{init}}=w_{c}$ a treatment by a reachability analysis is needed, as seen in the following Example 4.1.
{exa}
[…]
Figure 9. Example MDP $\mathcal{M}_{[c]}$
[…]
Figure 10. Transformed MDP $\mathcal{M}^{\max}_{[c]}$
@@
First, we show the soundness of Algorithm 4.1. By the virtue of Lemma 10 stating the soundness of the tranformation $\mathcal{M}$ to $\mathcal{M}_{[c]}$ it suffices to show that Algorithm 4.1 returns the correct answers “yes” or “no” when the task is to check whether the singleton $\mathsf{Cause}=\{c\}$ is an SPR cause in $\mathcal{N}=\mathcal{M}_{[c]}$. Recall the notation $q_{s}=\mathrm{Pr}^{\max}_{\mathcal{M}_{[c]},s}(\Diamond\mathsf{Eff})$. We abbreviate $q=q_{\mathsf{init}}$. Note that $(\neg\mathsf{Cause})\,\mathrm{U}\,c$ is equivalent to $\Diamond c$ as $c\in\mathsf{Cause}$.
@@
Obviously, statement (c) implies statements (a) and (b). The implication “(a) $\Longrightarrow$ (b)” follows from Lemma 5. We now turn to the proof of “(b) $\Longrightarrow$ (c)”. For this, we assume that we are given a GPR cause $\mathsf{Cause}$ for $\mathsf{Eff}$ in $\mathcal{M}$. For $c\in\mathsf{Cause}$, let $w_{c}=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$. Pick a state $c_{0}\in\mathsf{Cause}$ such that $w_{c_{0}}=\max\{w_{c}:c\in\mathsf{Cause}\}$. For every scheduler $\mathfrak{S}$ for $\mathcal{M}$ that minimizes the effect probability whenever it visits a state in $\mathsf{Cause}$, and visits $\mathsf{Cause}$ with positive probability, the conditional probability $\mathrm{Pr}^{\mathfrak{S}}_{\mathcal{M}}(\Diamond\mathsf{Eff}|\Diamond\mathsf{Cause})$ is a weighted average of the values $w_{c}$, $c\in\mathsf{Cause}$, and thus bounded by $w_{c_{0}}$. Using Lemma 10 we see that it is sufficient to only consider the minimal probabilities $w_{c}=\mathrm{Pr}^{\min}_{\mathcal{M},c}(\Diamond\mathsf{Eff})$. Thus, we conclude that $\{c_{0}\}$ is both an SPR and a GPR cause for $\mathsf{Eff}$.
@@ 4.2. Checking the global probability-raising condition
The remainder of this section is concerned with the proofs of both Theorem 17 and Theorem 18. For this, we suppose that $\mathsf{Cause}$ satisfies for all $c\in\mathsf{Cause}:\mathrm{Pr}^{\max}_{\mathcal{M}}(\neg\mathsf{Cause}\,\mathrm{U}\,c)>0$ which can be checked preemptively in polynomial time as argued before.
@@
We will start with a polynomial-time model transformation into a kind of “canonical form” after which we can make the following assumptions when checking the GPR condition of $\mathsf{Cause}$ for $\mathsf{Eff}$
@@
For every $c\in\mathsf{Cause}$, there is a single enabled action $\mathit{Act}(c)=\{\gamma\}$, and there is $w_{c}\in[0,1]\cap\mathbb{Q}$ such that $P(c,\gamma,\mathsf{eff}_{\mathsf{cov}})=w_{c}$ and $P(c,\gamma,\mathsf{noeff}_{\mathsf{fp}})=1{-}w_{c}$, where $\mathsf{noeff}_{\mathsf{fp}}$ is a terminal non-effect state and $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{eff}_{\mathsf{cov}}$ are only accessible via $\gamma$-transition from the $c\in\mathsf{Cause}$.
@@
The terminal states $\mathsf{eff}_{\mathsf{unc}}$, $\mathsf{eff}_{\mathsf{cov}}$, $\mathsf{noeff}_{\mathsf{fp}}$ and $\mathsf{noeff}_{\mathsf{tn}}$ are pairwise distinct. $\mathcal{M}$ can have further terminal states representing true negatives. These can be identified with $\mathsf{noeff}_{\mathsf{tn}}$.
@@
We justify the assumptions as we can transform $\mathcal{M}$ into a new MDP of the same asymptotic size satisfying the above assumptions. Thanks to Lemma 10, we may suppose that $\mathcal{M}=\mathcal{M}_{[\mathsf{Cause}]}$ without changing the satisfaction of (G). Thus, from cause states $c\in\mathsf{Cause}$ there are only two outgoing transitions, either to a terminal effect state $\mathsf{eff}$ with probability $\mathrm{Pr}^{\min}_{c}(\lozenge\mathsf{Eff})$ or to a terminal non-effect state $\mathsf{noeff}$ with the remaining probability (see Notation 4). We then may rename the effect state $\mathsf{eff}$ and the non-effect state $\mathsf{noeff}$ reachable from $\mathsf{Cause}$ into $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$, respectively. Furthermore, we collapse all other effect states into a single state $\mathsf{eff}_{\mathsf{unc}}$ and all true negative states into $\mathsf{noeff}_{\mathsf{tn}}$. Similarly, by renaming and possibly duplicating terminal states we also suppose that $\mathsf{noeff}_{\mathsf{fp}}$ has no other incoming transitions than the $\gamma$-transitions from the states in $\mathsf{Cause}$. This ensures (A1) and (A2). For (A3) consider the set $T$ of terminal states in the MDP obtained so far. We remove all non-trivial end components by switching to the MEC-quotient [dA97], i.e., we collapse all states that belong to the same MEC $\mathcal{E}$ into a single state $s_{\mathcal{E}}$ while ignoring the actions inside $\mathcal{E}$. Additionally, we add a fresh $\tau$-transition from the states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ (i.e., $P(s_{\mathcal{E}},\tau,\mathsf{noeff}_{\mathsf{tn}})=1$). The $\tau$-transitions from states $s_{\mathcal{E}}$ to $\mathsf{noeff}_{\mathsf{tn}}$ mimic cases where the scheduler of the original MDP enters the end component $\mathcal{E}$ and stays there forever.
@@
{exa}
For a demonstration of the above described transformations consider the abstract MDP $\mathcal{M}$ from Figure 12, where the dotted circles correspond to sets of states in the MDP. The MDP already has the form $\mathcal{M}=\mathcal{M}_{[c]}$. Now rename the effect state $\mathsf{eff}$ reachable from $\mathsf{Cause}$ to $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}$ to $\mathsf{noeff}_{\mathsf{fp}}$. All effect states which are not reachable from $\mathsf{Cause}$ are collapsed into the state $\mathsf{eff}_{\mathsf{unc}}$. In this example there are no terminal non-effect states which are not reachable from $c$. These would be collapsed to $\mathsf{noeff}_{\mathsf{tn}}$. Taking the MEC quotient collapses the MECs to states $s_{\mathcal{E}_{i}}$ while keeping outgoing transitions. Additionally, there is a fresh outgoing action $\tau$ from the states $s_{\mathcal{E}_{i}}$ to the fresh state $\mathsf{noeff}_{\mathsf{tn}}$. Thus we get $\mathcal{N}$ from Figure 12.
[…]
Figure 11. Example MDP $\mathcal{M}$ not satisfying
assumptions (A1)-(A3)
[…]
Figure 12. Transformed MDP $\mathcal{N}$ satisfying
assumptions (A1)-(A3)
[…]
The following Lemma 19 and Corollary 20 prove the soundness of the model transformation.
@@
By Lemma 3, there is a scheduler $\mathfrak{T}$ for $\mathcal{N}$ for each scheduler $\mathfrak{S}$ for $\mathcal{M}_{[\mathsf{Cause}]}$ such that each terminal state is reached with the same probability under $\mathfrak{T}$ in $\mathcal{N}$ and under $\mathfrak{S}$ in $\mathcal{M}_{[\mathsf{Cause}]}$. The state $\mathsf{eff}_{\mathsf{cov}}$ is present in $\mathcal{M}_{[\mathsf{Cause}]}$ under the name $\mathsf{eff}$. The state $\mathsf{eff}$ is furthermore reached in $\mathcal{M}_{[\mathsf{Cause}]}$ if and only if $\Diamond\mathsf{Cause}\land\Diamond\mathsf{Eff}$ is satisfied along a run. The set of terminal states in $\mathsf{Eff}$ is obtained from the set $\mathsf{Eff}$ in $\mathcal{M}_{[\mathsf{Cause}]}$ by collapsing states. As a scheduler $\mathfrak{S}$ can be viewed as a scheduler for both MDPs and these MDPs agree except for the terminal states, the first equality follows as well. As the probability to reach $\mathsf{Cause}$ is the sum of the probabilities to reach the terminal states $\mathsf{eff}_{\mathsf{cov}}$ and $\mathsf{noeff}_{\mathsf{fp}}$ in $\mathcal{N}$ and as these states are only renamed in $\mathcal{N}$ in comparison to $\mathcal{M}_{[\mathsf{Cause}]}$, the claim follows.
@@
Note, however, that the transformation changes the memory-requirements of schedulers witnessing that $\mathsf{Cause}$ is not a GPR cause for $\mathsf{Eff}$. We will address the memory requirements in the original MDP later. With assumptions (A1)-(A3), condition (G) can be reformulated as follows:
@@
With assumptions (A1)-(A3), a terminal state of $\mathcal{M}$ is reached almost surely under any scheduler after finitely many steps in expectation. Given a scheduler $\mathfrak{S}$ for $\mathcal{M}$ recall the definition of expected frequencies of state action-pairs $(s,\alpha)$, states $s\in S$ and state-sets $T\subseteq S$ under $\mathfrak{S}$:
@@
We can now prove that deciding the GPR condition can be done in $\mathrm{coNP}$, Theorem 17.
@@
Let $B_{2}$ be an open ball in $\mathbb{R}^{r}$ such that $h(B_{2})\subseteq B_{1}$ and such that $h(B_{2})$ contains $(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. We claim that $g\circ h\colon B_{2}\to\mathbb{R}$ has an isolated local maximum at $z^{\ast}\stackrel{{\scriptstyle\text{\rm\tiny def}}}{{=}}h^{-1}(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. It is clear that $g\circ h$ has a local maximum since $g$ has a local maximum at $(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. Suppose now, that $g\circ h$ does not have an isolated local maximum at $h^{-1}(y^{\ast},y_{\mathsf{Cause}}^{\ast})$. As $h$ is an affine map and the graph of $g$ is the solution to a quadratic equation, this is only possible if there is a direction $d\in\mathbb{R}^{r}\setminus\{0\}$ such that